# Gaudin models and moduli space of flower curves

### v1

@@ 5.2. Classical trigonometric Gaudin algebras
It particular, it is independent of $\theta$ and hence $A^{trig}_{\theta}(z_{1},\dots,z_{n})$ is also independent of $\theta$.
@@
The algebra $A^{trig}_{\theta}(z_{1},\dots,z_{n})$ is a polynomial algebra with generators $\Phi^{trig,m}_{l,i}$ for $l=1,\dots,r$, $i=1,\dots,n$, and $m=1,\dots d_{l}$.
@@
So it suffices to check that the generators $\{\Phi^{trig,m}_{l,i}\}$ for $l=1,\dots,r$, $i=1,\dots,n$, and $m=1,\dots d_{l}$ are algebraically independent.
@@
We will show that $V$ contains $\mathfrak{b}_{-}\oplus 0$ and that the image of $V$ in $\mathfrak{g}^{\oplus n-1}$ (projecting to the last $n-1$ summands) is $\mathfrak{b}_{-}^{\oplus n-1}$. Thus $\dim V=n\dim\mathfrak{b}_{-}$ as desired.
@@
The subalgebra generated by such coefficients coincides with the subalgebra spanned by all $p_{l,a}:=\Phi_{l}(x^{(1)}+a\pi_{-}(x^{(1)}))\in\mathcal{O}(\mathfrak{g}^{\oplus n})$ for $a\in\mathbb{C}$. This from the fact that a rational function with one pole is uniquely determined by the principal part of its Laurent expansion at this pole. We can regard $p_{l,a}$ as the composition of $\Phi_{l}$ with the linear map $\mathfrak{g}^{\oplus n}\rightarrow\mathfrak{g}$ given by $(x_{1},\dots,x_{n})\mapsto x_{1}+a\pi_{-}(x_{1})$.
@@ 5.3. Definition from representations
Let $M(\theta)=U\mathfrak{g}\otimes_{U\mathfrak{b}}\mathbb{C}v_{\theta}$ be the Verma module with highest weight $\theta$ (once again, we are identifying $\mathfrak{h}$ and $\mathfrak{h}^{*}$). Let $\alpha_{\theta}:M(\theta)\rightarrow\mathbb{C}$ be the linear map sending $v_{\theta}$ to 1 and sending other weight vectors to 0.
@@
For generic $\theta$, we know that $M(\theta)\otimes V\cong\oplus_{\nu}M(\theta+\nu)\otimes V_{\nu}$, where the direct sum is over the weights of $V$. Also for generic $\theta$, each $M(\theta+\nu)$ is irreducible and so $\operatorname{Hom}(M(\theta+\mu),M(\theta+\nu))$ is 0 unless $\mu=\nu$, in which case it is one dimensional. This establishes that $\operatorname{Hom}_{\mathfrak{g}}(M(\theta+\mu),M(\theta)\otimes V)$ and $V_{\mu}$ have the same dimension.
On the other hand, suppose that $\varphi\in\operatorname{Hom}(M(\theta+\mu),M(\theta)\otimes V)$ and $(\alpha_{\theta}\otimes\operatorname{id})(\varphi(v_{\theta+\mu}))=0$. Then $\varphi(v_{\theta+\mu})\in M(\theta)\otimes V$ is a singular vector (annihilated by $\mathfrak{n}_{+}$) and so the maximal non-zero weight component of its projection to $M(\theta)$ is also a singular vector. On the other hand, as $\theta$ is generic, $M(\theta)$ has no singular vector of weight less than $\theta$. Thus $\varphi=0$. ∎
@@
Let $\theta$ be integral and sufficiently dominant. Then the conclusion of Proposition 5.10 also holds if $M(\theta+\mu)$ and $M(\theta)$ are replaced by $V(\theta+\mu)$ and $V(\theta)$.
Consider the action of $\mathcal{A}(0,z_{1},\ldots,z_{n})$ on $M(\theta)\otimes V({\underline{\lambda}})$. Since the Gaudin algebra commutes with the diagonal $\mathfrak{g}$ action, there is an algebra homomorphism $\mathcal{A}(0,z_{1},\dots,z_{n})\rightarrow\operatorname{End}_{\mathfrak{g}}(M(\theta)\otimes V({\underline{\lambda}}))$ and thus $\mathcal{A}(0,z_{1},\dots,z_{n})$ acts on $\operatorname{Hom}_{\mathfrak{g}}(M(\theta+\mu),M(\theta)\otimes V({\underline{\lambda}}))$ for any $\mu$. Via Proposition 5.10, we have an isomorphism $\operatorname{Hom}_{\mathfrak{g}}(M(\theta+\mu),M(\theta)\otimes V({\underline{\lambda}}))\cong V({\underline{\lambda}})_{\mu}$ and thus an action of $\mathcal{A}(0,z_{1},\dots,z_{n})$ on $V({\underline{\lambda}})_{\mu}$.
@@
[…]
@@
The trigonometric Gaudin subalgebras can be described universally. Namely, there are two versions of a universal trigonometric Gaudin subalgebra, $\mathcal{A}^{trig}_{\theta}\subset U(\hat{\mathfrak{n}}_{+})$ and $\widetilde{\mathcal{A}}^{trig}_{\theta}\subset U(\hat{\mathfrak{b}}_{-})$ where $\hat{\mathfrak{n}}_{+}:=\mathfrak{n}_{+}+t^{-1}\mathfrak{g}[t^{-1}]$ and $\hat{\mathfrak{b}}_{-}:=\mathfrak{b}_{-}+t^{-1}\mathfrak{g}[t^{-1}]$.
@@
The subalgebra $U(\mathfrak{g}[t^{-1}])^{\mathfrak{g}}$ of $\mathfrak{g}$-invariants in $U(\mathfrak{g}[t^{-1}])$ naturally maps to this Hamiltonian reduction. By composition, we have a homomorphism $U(\mathfrak{g}[t^{-1}])^{\mathfrak{g}}\to U(\hat{\mathfrak{n}}_{+})$ and we let $\mathcal{A}_{\theta}^{trig}\subset U(\hat{\mathfrak{n}}_{+})$ be the image of $\widetilde{\mathcal{A}}\subset U(\mathfrak{g}[t^{-1}])^{\mathfrak{g}}$ under this homomorphism.
@@
Again by composition, we get a homomorphism $U(\mathfrak{g}[t^{-1}])^{\mathfrak{g}}\to U(\hat{\mathfrak{b}}_{-})$. We let $\widetilde{\mathcal{A}}^{trig}\subset U(\hat{\mathfrak{b}})$ be the image of $\widetilde{\mathcal{A}}\subset U(\mathfrak{g}[t^{-1}])^{\mathfrak{g}}$ under the homomorphism $i$; it is naturally a commutative subalgebra.
@@
$\widetilde{\mathcal{A}}^{trig}_{\theta}:=(ev_{\infty}^{\theta}\otimes\operatorname{id})\circ\Delta(\widetilde{\mathcal{A}}^{trig})\subset U(\hat{\mathfrak{b}}_{-}).$
@@
where the first arrow is the natural projection to the quantum Hamiltonian reduction by the diagonal $\mathfrak{g}$, the second one embeds it to the universal enveloping of the complement subalgebra $\mathfrak{n}_{+}\oplus\mathfrak{b}_{-}\oplus\mathfrak{g}^{\oplus n}$, and the last arrow takes the 00 character on $\mathfrak{n}_{+}$ and $\theta$ on $\mathfrak{b}_{-}$.
@@
    The subalgebra $\widetilde{\mathcal{A}}^{trig}_{\theta}\subset U(\hat{\mathfrak{b}}_{-})$ contains the quadratic $\widetilde{\Gamma}_{\theta}^{trig}:=\theta[-1]+\sum\limits_{a}h_{a}[-1]h^{a}[0]+\sum\limits_{a}e_{a}[-1]f_{a}[0]$. Moreover, the centralizer of $\widetilde{\Gamma}_{\theta}^{trig}$ in $U(\hat{\mathfrak{b}}_{-})$ is generated by $\widetilde{\mathcal{A}}^{trig}_{\theta}$ and $\mathfrak{h}[0]$.
@@
    The subalgebra $\mathcal{A}^{trig}_{\theta}\subset U(\hat{\mathfrak{n}}_{+})$ contains the quadratic $\Gamma_{\theta}^{trig}:=\theta[-1]+\sum\limits_{a}f_{a}[-1]e_{a}[0]$. Moreover $\mathcal{A}^{trig}_{\theta}$ is the centralizer of $\Gamma_{\theta}^{trig}$ in $U(\hat{\mathfrak{n}}_{+})$.
@@
    We have to show that the centralizer is not bigger than expected. So it is sufficient to see it in the classical limit, i.e. we have to check that the Poisson centralizers of the leading terms (with respect to the PBW filtration) of $\widetilde{\Gamma}_{\theta}^{trig}$, $\Gamma_{\theta}^{trig}$ in $S\hat{\mathfrak{b}}$, $S\hat{\mathfrak{n}}$ are $\operatorname{gr}\widetilde{\mathcal{A}}^{trig}_{\theta}\cdot S\mathfrak{h}[0]$ and $\mathcal{A}^{trig}_{\theta}$, respectively. Moreover, it suffices to check that the Poincaré series of the centralizers in question are not greater than those expected.
@@
    There is a $1$-parametric family $\widetilde{\varphi}_{s}$ of Poisson automorphisms of $S\hat{\mathfrak{b}}$ such that
    $\lim\limits_{s\to\infty}s^{-1}\widetilde{\varphi}_{s}(\widetilde{\Gamma}^{trig}_{\theta})=\theta[-1]\text{ and }\lim\limits_{s\to\infty}\widetilde{\varphi}_{s}(\operatorname{gr}\widetilde{\mathcal{A}}^{trig}_{\theta}\cdot S\mathfrak{h}[0])=S\mathfrak{h}[t^{-1}].$
    Namely, we define $\widetilde{\varphi}_{s}(h_{a}[0]):=s(\theta,h_{a})$ and $\widetilde{\varphi}_{s}(e_{a}[0])=\widetilde{\varphi}_{s}(e_{a}[k])=\widetilde{\varphi}_{s}(f_{a}[k])=\widetilde{\varphi}_{s}(h_{a}[k])=0$ for all $k<0$.
    Similarly, the family $\varphi_{s}$ of Poisson automorphisms of $S\hat{\mathfrak{n}}$ defined as $\varphi_{s}(e_{a}[-1]):=s(f,e_{a})$ and $\varphi_{s}(e_{a}[k])=\varphi_{s}(f_{a}[k])=\varphi_{s}(h_{a}[k])=0$ for all $k<0$, with $f\in\mathfrak{g}$ being the principal nilpotent element, satisfies the properties $\lim\limits_{s\to\infty}s^{-1}\varphi_{s}(\Gamma^{trig}_{\theta})=f[-1]$ and $\lim\limits_{s\to\infty}\varphi_{s}(\operatorname{gr}\mathcal{A}^{trig}_{\theta})=S(t^{-1}\mathfrak{z}_{\mathfrak{g}}(f)[t^{-1}])$.
@@
    Note that $S\mathfrak{h}[t^{-1}]$ is exactly the Poisson centralizer of $\theta[-1]$ in $S\hat{\mathfrak{b}}$ and $S(t^{-1}\mathfrak{z}_{\mathfrak{g}}(f)[t^{-1}])$ is exactly the Poisson centralizer of $f[-1]$ in $S\hat{\mathfrak{n}}$, so the Poincare series of the centralizers are not bigger than the expected.
Now we can compare our trigonometric Gaudin subalgebras with those introduced by Molev and Ragoucy in [MR] in type $A$. Namely, in [MR], some higher Hamiltonians in $U(\hat{\mathfrak{b}}_{+})$ are constructed (i.e. the roles of positive and negative roots are switched, compared to ours).
@@
The higher trigonometric Gaudin Hamiltonians in type A from [MR, Section 2] lie in $\tilde{\mathcal{A}}^{trig}_{0}\subset U(\hat{\mathfrak{b}}_{+})$.
@@
The higher Hamiltonians described in [MR] come from a universal commutative subalgebra in $U(\hat{\mathfrak{b}}_{+})$ that contains $\Gamma_{0}^{trig}$. ∎

### v2

@@ 5.2. Classical trigonometric Gaudin algebras
In particular, it is independent of $\theta$ and hence $A^{trig}_{\theta}(z_{1},\dots,z_{n})$ is also independent of $\theta$.
@@
The algebra $A^{trig}_{\theta}(z_{1},\dots,z_{n})$ is a polynomial algebra with generators $\Phi^{trig,m}_{l,i}$ for $l=1,\dots,r$, $i=1,\dots,n$, and $m=1,\dots,d_{l}$.
@@
So it suffices to check that the generators $\{\Phi^{trig,m}_{l,i}\}$ for $l=1,\dots,r$, $i=1,\dots,n$, and $m=1,\dots,d_{l}$ are algebraically independent.
@@
We will show that $V$ contains $\mathfrak{b}_{-}\oplus 0$ and that the image of $V$ in $\mathfrak{g}^{\oplus n-1}$ (projecting to the last n1n-1 summands) is $\mathfrak{b}_{-}^{\oplus n-1}$. Thus $\dim V=n\dim\mathfrak{b}_{-}$ as desired.
@@
The subalgebra generated by such coefficients coincides with the subalgebra spanned by all $p_{l,a}:=\Phi_{l}(x^{(1)}+a\pi_{-}(x^{(1)}))\in\mathcal{O}(\mathfrak{g}^{\oplus n})$ for $a\in\mathbb{C}$. This follows from the fact that a rational function with one pole is uniquely determined by the principal part of its Laurent expansion at this pole. We can regard $p_{l,a}$ as the composition of $\Phi_{l}$ with the linear map $\mathfrak{g}^{\oplus n}\rightarrow\mathfrak{g}$ given by $(x_{1},\dots,x_{n})\mapsto x_{1}+a\pi_{-}(x_{1})$.
@@ 5.3. Definition from representations
Let $M(\theta)=U\mathfrak{g}\otimes_{U\mathfrak{b}}\mathbb{C}v_{\theta}$ be the Verma module with highest weight $\theta$ (once again, we are identifying $\mathfrak{h}$ and $\mathfrak{h}^{*}$). We denote by $M^{\vee}(\theta)$ the contragredient dual of $M(\theta)$ and by $s_{\theta}:M(\theta)\to M^{\vee}(\theta)$ the (unique up to proportionality) homomorphism given by the Shapovalov pairing. $M^{\vee}(\theta)$ has the same character as $M(\theta)$ and is cofree with $1$ co-generator as a $U(\mathfrak{n}_{+})$-module. Let $\alpha^{\vee}_{\theta}:M^{\vee}(\theta)\rightarrow\mathbb{C}$ be the (unique up to proportionality) nonzero linear map sending all the weight vectors with the weights different from $\theta$ to 0. Similarly, let $\alpha_{\theta}=\alpha^{\vee}_{\theta}\circ s_{\theta}:M(\theta)\rightarrow\mathbb{C}$ be the linear map sending $v_{\theta}$ to 1 and all the weight vectors of other weights to 0.
@@
 [^4] We thank the referee for pointing out the natural generality of this statement and for suggesting this argument.
Let $V$ be any finite-dimensional representation of $\mathfrak{g}$ and let $\mu$ be an integral weight. Then the linear map
$\pi^{\vee}_{\theta}:\operatorname{Hom}_{\mathfrak{g}}(M(\theta+\mu),M^{\vee}(\theta)\otimes V)\rightarrow V_{\mu}$
given by $\varphi\mapsto(\alpha^{\vee}_{\theta}\otimes\operatorname{id})(\varphi(v_{\theta+\mu}))$ is an isomorphism.
Proof.
Since $M(\theta)$ is a $\mathfrak{g}$-module induced from the $1$-dimensional $\mathfrak{b}$-module $\mathbb{C}v_{\theta}$, we have, for any $\mathfrak{g}$-module $L$, $\operatorname{Hom}_{\mathfrak{g}}(M(\theta),L)=\operatorname{Hom}_{\mathfrak{b}}(\mathbb{C}v_{\theta},L)$. Next, $M^{\vee}(\theta)$ is $U(\mathfrak{n}_{+})$-cofree with one generator of the weight $\theta$, i.e. it is the graded dual $\mathfrak{b}$-module of the induced from the $1$-dimensional $\mathfrak{h}$-module $\mathbb{C}v_{\theta}$. So, for any finite-dimensional $\mathfrak{b}$-modules $K,L$ that are semisimple with respect to $\mathfrak{h}$, we have $\operatorname{Hom}_{\mathfrak{b}}(K,M^{\vee}(\theta)\otimes L)=\operatorname{Hom}_{\mathfrak{b}}(K\otimes\mathbb{C}v_{\theta}^{*},L)$. So we get
$\operatorname{Hom}_{\mathfrak{g}}(M(\theta+\mu),M^{\vee}(\theta)\otimes V)=\operatorname{Hom}_{\mathfrak{b}}(\mathbb{C}v_{\theta+\mu},M^{\vee}(\theta)\otimes V)=\operatorname{Hom}_{\mathfrak{h}}(\mathbb{C}v_{\theta+\mu}\otimes\mathbb{C}v_{\theta}^{*},V)=V_{\mu}.$
This establishes that $\operatorname{Hom}_{\mathfrak{g}}(M(\theta+\mu),M^{\vee}(\theta)\otimes V)$ and $V_{\mu}$ have the same dimension. On the other hand, suppose that $\varphi\in\operatorname{Hom}(M(\theta+\mu),M(\theta)\otimes V)$ and $(\alpha^{\vee}_{\theta}\otimes\operatorname{id})(\varphi(v_{\theta+\mu}))=0$. Then $\varphi(v_{\theta+\mu})\in M^{\vee}(\theta)\otimes V$ is a singular vector (annihilated by $\mathfrak{n}_{+}$) and so the maximal non-zero weight component of its projection to $M^{\vee}(\theta)$ is also a singular vector. As $M^{\vee}(\theta)$ is $U(\mathfrak{n}_{+})$-cofree, it has no singular vector of weight less than $\theta$. Thus $\varphi=0$.
Corollary 5.11.
@@
This follows from Proposition 4 since for generic $\theta$ we have $M(\theta)=M^{\vee}(\theta)$.
An alternate (and a bit more elementary) proof is as follows. For generic $\theta$ we have $M(\theta)\otimes V\cong\oplus_{\nu}M(\theta+\nu)\otimes V_{\nu}$, where the direct sum is over the weights of $V$. Also for generic $\theta$, each $M(\theta+\nu)$ is irreducible and so $\operatorname{Hom}(M(\theta+\mu),M(\theta+\nu))$ is 0 unless $\mu=\nu$, in which case it is one dimensional. This shows that $\operatorname{Hom}_{\mathfrak{g}}(M(\theta+\mu),M(\theta)\otimes V)$ and $V_{\mu}$ have the same dimension. The argument proving that $\pi_{\theta}$ is an isomorphism is the same as in Proposition 4. ∎
@@
Let $\theta$ be integral and sufficiently dominant. Then the statement of Proposition 4 holds true if $M(\theta+\mu)$ and $M(\theta)$ are replaced by $V(\theta+\mu)$ and $V(\theta)$.
Consider the action of $\mathcal{A}(0,z_{1},\ldots,z_{n})$ on $M(\theta)\otimes V({\underline{\lambda}})$. Since the Gaudin algebra commutes with the diagonal $\mathfrak{g}$ action, there is an algebra homomorphism $\mathcal{A}(0,z_{1},\dots,z_{n})\rightarrow\operatorname{End}_{\mathfrak{g}}(M(\theta)\otimes V({\underline{\lambda}}))$ and thus $\mathcal{A}(0,z_{1},\dots,z_{n})$ acts on $\operatorname{Hom}_{\mathfrak{g}}(M(\theta+\mu),M(\theta)\otimes V({\underline{\lambda}}))$ for any $\mu$. Via Proposition 4, we have an isomorphism $\operatorname{Hom}_{\mathfrak{g}}(M(\theta+\mu),M(\theta)\otimes V({\underline{\lambda}}))\cong V({\underline{\lambda}})_{\mu}$ and thus an action of $\mathcal{A}(0,z_{1},\dots,z_{n})$ on $V({\underline{\lambda}})_{\mu}$.
@@
Now, if $u\in M(\theta)$ and $y\in U\mathfrak{b}_{-}$, then $\alpha_{\theta}(y\cdot u)=\theta(y)\alpha_{\theta}(u)$, where by a slight abuse of notation, we use $\theta$ for the composition $U\mathfrak{b}_{-}\rightarrow U\mathfrak{h}\xrightarrow{\theta}\mathbb{C}$. Thus since $x^{\prime\prime}\in U\mathfrak{b}_{-}\otimes U\mathfrak{g}^{\otimes n}$, we have $(\alpha_{\theta}\otimes id)(x^{\prime\prime}\cdot v)=(\theta\otimes id)(x^{\prime\prime})\cdot(\alpha_{\theta}\otimes id)(v)$.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
where the first arrow is the natural projection to the quantum Hamiltonian reduction by the diagonal $\mathfrak{g}$, the second one embeds it to the universal enveloping of the complement subalgebra $\mathfrak{n}_{+}\oplus\mathfrak{b}_{-}\oplus\mathfrak{g}^{\oplus n}$, and the last arrow takes the 0 character on $\mathfrak{n}_{+}$ and $\theta$ on $\mathfrak{b}_{-}$.
@@
[…]
@@
[…]
@@
    We have to show that the centralizer is not bigger than expected. So it is sufficient to see it in the classical limit, i.e. we have to check that the Poisson centralizers of the leading terms (with respect to the PBW filtration) of $\widetilde{\Gamma}_{\theta}^{trig}$, $\Gamma_{\theta}^{trig}$ in $S\hat{\mathfrak{b}}$, $S\hat{\mathfrak{n}}$ are $\operatorname{gr}\widetilde{\mathcal{A}}^{trig}_{\theta}\cdot S\mathfrak{h}[0]$ and $\mathcal{A}^{trig}_{\theta}$, respectively. Moreover, it suffices to check that the Hilbert-Poincaré series of the centralizers in question are not greater than those expected. Slightly abusing notations, we denote the leading terms of $\widetilde{\Gamma}_{\theta}^{trig}$, $\Gamma_{\theta}^{trig}$ in $S\hat{\mathfrak{b}}$, $S\hat{\mathfrak{n}}$ by the same letters $\widetilde{\Gamma}_{\theta}^{trig}$ and $\Gamma_{\theta}^{trig}$.
@@
    To any regular $\eta\in\mathfrak{h}$, we assign a $1$-parametric family $\widetilde{\varphi}_{s}$ of Poisson automorphisms of $S\hat{\mathfrak{b}}$ such that
    $\lim\limits_{s\to\infty}s^{-1}\widetilde{\varphi}_{s}(\widetilde{\Gamma}^{trig}_{\theta})=\eta[-1]\text{ and }\lim\limits_{s\to\infty}\widetilde{\varphi}_{s}(\operatorname{gr}\widetilde{\mathcal{A}}^{trig}_{\theta}\cdot S\mathfrak{h}[0])=S\mathfrak{h}[t^{-1}].$
    Namely, we define $\widetilde{\varphi}_{s}(h_{a}[0]):=h_{a}[0]+s(\eta,h_{a})$ and $\widetilde{\varphi}_{s}(e_{a}[0])=e_{a}[0];\ \widetilde{\varphi}_{s}(e_{a}[k])=e_{a}[k];\ \widetilde{\varphi}_{s}(f_{a}[k])=f_{a}[k];\ \widetilde{\varphi}_{s}(h_{a}[k])=h_{a}[k]$ for all $k<0$.
    Similarly, to the principal nilpotent element $f\in\mathfrak{n}_{-}\subset\mathfrak{g}$, we assign the family $\varphi_{s}$ of Poisson automorphisms of $S\hat{\mathfrak{n}}$ defined as $\varphi_{s}(e_{a}[0]):=e_{a}[0]+s(f,e_{a})$ and $\varphi_{s}(e_{a}[k])=e_{a}[k];\ \varphi_{s}(f_{a}[k])=f_{a}[k];\ \varphi_{s}(h_{a}[k])=h_{a}[k]$ for all $k<0$. This family of automorphisms satisfies the properties $\lim\limits_{s\to\infty}s^{-1}\varphi_{s}(\Gamma^{trig}_{\theta})=e[-1]$ and $\lim\limits_{s\to\infty}\varphi_{s}(\operatorname{gr}\mathcal{A}^{trig}_{\theta})=S(t^{-1}\mathfrak{z}_{\mathfrak{g}}(e)[t^{-1}])$, where $e\in\mathfrak{n}_{+}\subset\mathfrak{g}$ is the opposite principal nilpotent element.
@@
    It is suffices to check that the above limit subalgebras are Poisson centralizers of the limits of the quadratic elements, i.e. of $\eta[-1]$ and e[1]e[-1], respectively. Indeed, this implies that $\widetilde{\varphi}_{s}(\operatorname{gr}\widetilde{\mathcal{A}}^{trig}_{\theta})$ is the Poisson centralizer $\widetilde{\varphi}_{s}(\widetilde{\Gamma_{\theta}^{trig}}) fo$r $s$ large enough, and hence for any $s\in\mathbb{C}$ since $\widetilde{\varphi}_{s}$ is a Poisson automorphism. Similarly, for $\operatorname{gr}\mathcal{A}_{\theta}^{trig}$.
  4. (4)
    By the argument entirely similar to [R2, Lemma 4], $S\mathfrak{h}[t^{-1}]$ is exactly the Poisson centralizer of $\eta[-1]$ in $S\hat{\mathfrak{b}}$ and $S(t^{-1}\mathfrak{z}_{\mathfrak{g}}(e)[t^{-1}])$ is exactly the Poisson centralizer of e[1]e[-1] in $S\hat{\mathfrak{n}}$, so the Hilbert-Poincaré series of the centralizers are not bigger than the expected.
    Let us reproduce this argument for completeness. Choose a basis $\{x_{a}\}$ in $\mathfrak{g}$ compatible with $\mathrm{ad}_{\eta}$ (respectively, with $\mathrm{ad}_{e}$), i.e. an eigenbasis for $\mathrm{ad}_{\eta}$ and a Jordan basis for $\mathrm{ad}_{e}$, respectively. Multiplying its elements by negative powers of $t$, we get bases in $\hat{\mathfrak{b}}_{-}$ (respectively, in $\hat{\mathfrak{n}}_{+}$). Take any ordering of the basis $\{x_{a}\}$ such that for $\eta,x_{a}]=0,\ [\eta,x_{b}]\neq 0$ (respectively, $e,x_{a}]=0,\ [e,x_{b}]\neq 0$) we have $x_{a}<x_{b}$. In the second case, we need additionally $e,x_{b}]<x_{b}$. Extend these orders to the bases $\{x_{a}[k]\}$ in such a way that $x_{a}[k]<x_{b}[m]$ if $x_{a}<x_{b}$ or if $x_{a}=x_{b}$ and $k>m$. This determines the lexicographic ordering on the monomials of $\{x_{a}[k]\}$ preserved by $\mathrm{ad}_{\eta[-1]}$ (respectively, by $\mathrm{ad}_{e[-1]}$). Now it remains to notice that for any element not from $S(\mathfrak{h}[t^{-1}])$ (respectively, not from $S(t^{-1}\mathfrak{z}_{\mathfrak{g}}(e)[t^{-1}])$), its leading term is mot annihilated by $\mathrm{ad}_{\eta[-1]}$ (respectively, by $\mathrm{ad}_{e[-1]}$).
[…]
@@
[…]
@@
[…]